DID YOU DO ANYTHING IN CLASS TODAY?

Richard L. Francis

Southeast Missouri State University

The question posed by the title we’ve heard many times before. Frequent also is its future counterpart, that is, “Are you going to do anything in class tomorrow?” The student who makes the inquiry is likely, though subconsciously, saying something to the thoughtful listener. Teachers, in occasional but trying tests of their enthusiasm, are apt to answer facetiously, “No, not much.”

Why are such questions asked? Several possibilities come to mind. In the casual words of the student these possibilities include:

1. “Missing a single day of class can’t hurt.”

2. “You can’t do that much in one session.”

3. “It’s easy to make up missed course work.”

4. “The textbook covers what the teacher did anyway.”

5. “I get so little out of class when I do go.”

6. “Others miss class and survive. So can I.”

7. “I can pass the course without doing all the assigned work.”

All of these responses are negative in tone. In fairness, it must be noted that many learners are conscientious about their course work. They simply make a poor word choice in asking the question, “Did you do anything in class today?” Such students may really be asking, “Did I miss a test or some other urgent learning activity for which I am immediately accountable?” The more tactful and diplomatic question does not always find expression.

Adding “Zeros” and Getting Something

A limited perception still persists to the effect that the classroom routine is not eventful nor of long term impact. In this assessment of things, highlights such as major tests or presentations deviate from the routine and thus count for “having done something in class.” Beyond these, student comments as the value of a given class session may be filled with mixed and confusing signals.

I’ve occasionally asked students, “What did you learn today?” More often than I wanted to hear, the reply was, “Nothing at all.” Nevertheless, all these zeros of achievement, ultimately when added together, yielded something positive. He or she had become a better reader or a more capable solver of mathematics problems. The student had paradoxically moved to a higher plateau of achievement. A perceived nothing had culminated in something.

Enthusiasm and Patience

In weaker moments, enthusiasm may be of such a character that students fail to sense the importance of a given day’s work. There is likely a routine that encourages the question of this article’s title. Such a routine takes the form in mathematics of “getting on with the problems.” Falling thus, into a HO–HUM category, it complicates the student’s acquisition of a fair perspective. Teacher and student enthusiasm manifests itself in diverse ways and goes far in dispelling negative feelings which otherwise can characterize a class session. (See [1].)

The instructor, of course, must be mindful of a widespread student tendency to minimize the importance of a single day’s work in the classroom. As a consequence of this tendency the student should be encouraged to reflect, late in the class session, on the hour’s activities and ask himself/herself the probing question, “What did I learn today?” An extreme measure perhaps would be the posting, in a conspicuous place, of a plaque with the now familiar words

WHAT DID I LEARN IN CLASS TODAY?

The answer to such a summarizing question may be surprising. Maybe it is “I learned how to differentiate implicitly,” or “I now have a deeper understanding, by repeated application, of the delta process.”

And then there’s patience! A sign of the times is that of instant everything, be it the coffee we drink, the mastery of a musical instrument, or miracle photography. The learning of mathematics hardly occurs in such a rushed manner. Achieving the goal of mathematical maturity is one of a challenging, patient kind. It takes time to get there. The hour hand of a clock appears motionless. Yet, in the course of the day, we realize the opposite. Even as mathematical growth, reflecting the work of the classroom moments, appears imperceptible or as nothing, still in time can we see a picture of remarkable progress.

So, what really happened in class today? More perhaps than what an impulsive, superficial judgment might suggest.

\underline{Reference}

1. R. L. Francis, “He Knows His Subject, But …,” Missouri Schools, March 1975, 22–25.

23